OpenStudy (anonymous):

my problem is 20 + 8(q-11)=(-12)

OpenStudy (nikato):

WELCOME TO OPENSTUDY!

OpenStudy (nikato):

the first thing u would do is distribute the 8 to q-11

OpenStudy (anonymous):

after I distributed I got 8q - 88. Is that correct?

OpenStudy (sleepyjess):

Yes

OpenStudy (sleepyjess):

So now what is your whole equation

OpenStudy (anonymous):

20 + 8q -88 = (-12)

OpenStudy (sleepyjess):

Correct... what do you think you would do next?

OpenStudy (nikato):

yes, now combine like terms so 20-88

OpenStudy (sleepyjess):

lol... okay, since nikato gave you that one, what is your equation now?

OpenStudy (anonymous):

I'm a bit confused

OpenStudy (sleepyjess):

And then what would your next step be?

OpenStudy (sleepyjess):

about what?

OpenStudy (anonymous):

the equation

OpenStudy (sleepyjess):

20 + 8q -88 = (-12)

OpenStudy (sleepyjess):

then you combine like terms

OpenStudy (sleepyjess):

If it helps to reorder, you can do that first... 8q+20-88 = (-12)

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

so my equation should be 8q+68=(-12)

OpenStudy (sleepyjess):

Almost... 20-88 would be -68

OpenStudy (sleepyjess):

So your equation would be 8q-68=(-12)

OpenStudy (anonymous):

okay, I was confused about that part. Thank you for helping. I can do the rest from here. :)

OpenStudy (sleepyjess):

Now you want to get the q by itself, which is going to take 2 steps.

OpenStudy (sleepyjess):

Do you know how to do the rest then?

OpenStudy (anonymous):

yes, I have to do two-step equation.

OpenStudy (sleepyjess):

Okay... what's your final answer?

OpenStudy (anonymous):

q = 7

OpenStudy (sleepyjess):

Yep :-)

OpenStudy (anonymous):

thank you so much for your help! :)

OpenStudy (sleepyjess):

I hope you didn't mind me doing it that way with you... I just try to explain it like I would want it explained.

OpenStudy (sleepyjess):

yw

OpenStudy (anonymous):

I didn't mind at all. That how I like it to explained too.